Operations Report – December 21st

SOL:5

Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None
Generator (hours run): 17:27 last night until around 8:00 Today started at 17:22

Solar— SOC 79%%
Diesel Reading – 77%
Propane Reading – 80%
Ethanol Free Gasoline – N/A gallons.

Water (Axillary tank) – 0 gallons
Water (static tank) – High Level, around 300 -gallons
Axillary to Static tank transfer– no
Gallons transferred: 0
Water in GreenHab – Medium level (65%)
Water (loft) – Static to Loft Pump used – yes
Water Meter: 01392989

Toilet tank emptied: YES

Deimos rover used: no, still not functional
Hours: 0
Beginning charge: –
Ending charge: –
Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours: 63
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Opportunity rover used: NO
Hours: 44.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Curiosity rover used: NO
Hours: 60
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Notes on rovers: Nothing to report.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station.

HabCar used and why, where? Not used.
CrewCar used and why, where? Not used.

General notes and comments: – We had one walk EVA for the selection of some places to perform one experiment.

Summary of internet: Nothing to report.

Summary of suits and radios: – The 2-piece suits number 3, 4 and 6 were checked today. Number 3 is working fine. Number 4 had a bit cutted the inside mechanical joint, so the air flow was not the best. We temporary fix it. Suit number 6 had one stocked fan. We followed the instructions of Capcom in order to fix it and it is working fine. Find attached pictures
– The 1-piece suit number 8, 9 and 10 are superficially damaged. Find attached pictures.

Summary of GreenHab operations: – NOthing to report.

Summary of ScienceDome operations: – Nohing to report.

Summary of RAM operations: We have been working some engineering and experiments here. All good.

Summary of any observatory issues: Musk Solar observatory – Nominal. RObotic Observatory – In use at 17:06

Summary of health and safety issues: The provided smoke detector has been installed.

Questions, concerns and requests to Mission Support: Find attached the photos of the suits while repairing and the one that we saw with one damage.

Thank you!!!